Find the square of the following numbers. 
93
Find the square of the following numbers. 
93

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(93)^2
= (90+3 )^2
= (90)^2+ (3)^2+ 2×90×3 [Since, (a+b)^2= a^2+b^2+2ab]
=8100+9+540
= 8649

Thanks